A given mass of gas has a volume of 7.5cm3 at 250k and 80Nm^-2. Calculate the pressure at which it will have a volume of 7.2 cm3 at 270k
1
Expert's answer
2019-11-18T06:28:05-0500

(P1*V1)/(P1*V1)=T1/T2


P2=(P1*V1*T2)/(T1*V2)=(80 N*m^-2*7.5*10^-6m^3*270K)/(250K*7.2*10^-6m^3)=90 N*m^-2